# Speed-Accuracy Trade-Off in the Simulation of Realistic Arterial Blood Flow: Effects of Geometric Simplifications and Non-Newtonian Models

### v1

@@ Abstract
This study investigates various aspects of blood flow in realistic arteries obtained from Computed Tomography (CT) scan data and its geometrically simplified variants. The study uses realistic inflow and outlet boundary conditions with Newtonian and non-Newtonian blood viscosity models, including the one recently developed by Apostolidis and Beris [28]. Similar studies exist to predict the progression of atherosclerosis, clots, and thrombus formation in arteries. However, these assume smooth flow geometries. In contrast, realistic arteries possess significant amounts of roughness, perturbations, and undulations. To understand their impact, studies are performed with two levels of geometry simplifications. Our results indicate that flow field predictions on these simplified geometries are visibly affected. The first level of simplification loses the local asymmetry but is able to reasonably approximate various parameters, capturing patterns with the correct magnitude. However, further simplification to a smooth one does not show vortex structures in the flow field. The flow recirculation zones near wall perturbations progressively weaken as the geometry is smoothed. On the other hand, no significant differences are observed between Newtonian and non-Newtonian models in this study. Our analysis also enables a speed-accuracy trade-off for future applications. The number of iterations to convergence becomes progressively lower as geometry is simplified, with nearly four times the difference obtained even with the first level of simplification. Further, the computational time for the Newtonian model is nearly 10 times smaller than the non-Newtonian counterpart.
@@ 1 Introduction
Cardiovascular diseases (CVD) remain one primary reason for human mortality [1] and have grown in importance after the COVID-19 pandemic. The progression of CVD is directly related to the geometric details of the arteries [2]. CVD is also known to aggravate other conditions as well [3]. Over the last few decades, greater computing power has enabled detailed computational studies on blood flow in the human arteries [4, 5]. Researchers have been interested in the link between atherosclerosis and fluid flow and related parameters, including the wall shear stress (WSS) [6, 7, 8, 9]. Increasingly, studies have used patient-specific geometries with the aim of contributing to medical applications [10]. Similar studies have been reported to predict the progression of atherosclerosis [11, 12], clots [13], and formation of thrombus [14]. Significant studies have investigated the effect of stenosis on blood flow [15, 16, 17]. However, in all these studies, the flow geometries have been assumed to be quite regular. In contrast, human arteries are far from regular and possess a significant amount of undulations in diameter as well as in the flow direction. The geometrical variations of arteries have been linked to the onset of atherosclerosis [18, 19, 20], thereby justifying its importance in the broader aspect of human health.
@@
Despite the apparent importance of irregularities in geometry, researchers have rarely focused on their blood flow. Yi et al. [21] performed simulations to study the impact of surface roughness on arterial stenosis. Shukla et al.[22] studied the effect of surface roughness on pulsatile flow in a tube with a periodic pattern of roughness in the form of sinusoids while Owen et al. [23] studied the effect of surface roughness on blood flow in a tube. To the best of our knowledge, these three studies come close to studying the impact of roughness from a biological perspective for blood flow analysis. Among them, the studies by Yi et al. [21] and Shukla et al. [22] use Newtonian fluid models. Owen et al. [23] used a variety of models, including Newtonian and non-Newtonian ones, to study the effect of an uneven geometry. However, the geometry and setup they created in the study can be argued to be far removed from reality. In a tube, they made a ”rough” patch with surface irregularities that extended along the length and compared fluid flow results between the rough patch and an equivalent smooth patch on the other side. This geometrical setup has two significant problems. Firstly, the roughness is distributed all over the surface, and a tube with a rough and smooth patch is not close to the surface features of arteries in our body. Secondly, the surface features comprise not just roughness but also changes on the scale of the artery diameter. Hence, irregularities also exist in the shape and size fluctuations along the axial length of the arterial tube. The flowing fluid experiences the combined effect of these irregularities. As discussed later, the real artery may deviate significantly from the circular shape over a given plane. Additionally, the equivalent diameter can also show variations along the axis. To the best of our knowledge, no previous study has comprehensively investigated the effect of such shape and size perturbations, in addition to surface roughness, which is aimed for in the present study. Also, we use a variety of viscosity models to examine the sensitivity of the non-Newtonian nature of blood with geometric perturbations.
@@
One additional aspect studied here is the change obtained by the choice of the viscosity model. Calculations in these three equivalent geometric representations provide us with intricate details of the effects of geometric perturbations on the flow of blood and their connection with the chosen rheological model, apart from the effect of these details on the computational speed.
The outcomes of the simulations are shown in terms of vortical patterns and hemodynamic parameters such as the time average wall shear stress (TAWSS) and oscillatory shear index (OSI). The differences in flow patterns among the three levels are discussed in detail, with an understanding of the reduction in CPU time, which is substantial. The results presented in this study may enable the simplification of complex arterial networks to increase the feasibility of meaningful full-body simulations in real-time.
@@ 2.1 Geometry
Figure 1: Level of geometrical aspects is shown, in (i) Level 1 of descending aorta extracted from CT-scan data, (i) Slice-based geometry of descending aorta (Level 1 is divided into 25 slices for the construction of level 2). (iii) Level 3 geometry based on the inlet and outlet diameter of Level 1 geometry. (iv) Shows the cross-section area variation over the normalized z-axis for each level of descending aorta simplification.
@@ 2.2 Governing equations
Patient-specific and simplified models of the arteries are used to study the hemodynamics of blood flow under pulsatile flow conditions. In this respect, the incompressible form of the continuity and momentum equations are solved with realistic boundary conditions with appropriate viscosity models. The incompressibility constraint in the form of the continuity equation and the applicable momentum equations for a variable viscosity fluid are given below.
@@ 2.2.1 Boundary Conditions
To solve the governing equations for the descending aorta and its simplified geometric versions, the following boundary conditions are applied:
@@
    Inlet conditions: A uniform inflow velocity profile is imposed at the inlets of all geometries, combined with a pulsatile waveform to replicate the periodic nature of blood flow under physiological conditions. This pulsatile waveform captures the transient characteristics of flow and ensures a realistic simulation of the hemodynamic environment. The pulsatile velocity profile is derived from physiological measurements reported by Mills et al [25]. The derived profile is fitted by a Fourier series of up to 18 terms to apply the inlet boundary condition. To fully allow the inflow to develop into a Womersley profile, the inlet region of the artery is extended by a length of $10D$, where $D$ represents the characteristic diameter of the artery. This extension minimizes upstream flow irregularities owing to flow development and ensures a realistic inflow condition in the computational domain of interest. Similarly, an extension of $5D$ is added to the outlet to mitigate numerical reflections of an outflow boundary condition and maintain a smooth flow transition at the exit plane.
@@
    Outlet conditions: Realistic pulsatile gauge pressure is applied at the outlets of the geometries, as depicted in Figure 2. This boundary condition mirrors the dynamic pressure variations observed in the cardiovascular system, which are critical for accurately capturing the pressure-driven nature of blood flow. The pulsatile pressure profile is derived from physiological measurements carried out by Mills et al [25], ensuring that they impose real-world conditions. These outlet conditions play a key role in maintaining the accuracy of flow characteristics, especially in scenarios where downstream vascular resistance and compliance significantly affect the hemodynamics.
@@
    Wall conditions: The arterial walls in all geometries are assumed to be rigid, i.e. they do not deform under the influence of pressure and shear loading. A no-slip boundary condition is enforced at the walls, requiring the velocity at the wall surface is zero. This assumption is widely used in computational blood flow simulations.
@@ 2.2.2 Blood viscosity models
    In this context, $\mu_{0}$ and $\mu_{\infty}$ are the saturation values of viscosity at low and high shear rates, respectively. The parameter $P$ is known as the Yasuda exponent, with both $P$ and $n$ determining the degree of nonlinearity. The time constant $\lambda$ regulates the range of strain rates at which the transition to a power-law type regime is obtained. In a way, $\lambda$ is related to the relaxation time of the fluid. The model simplifies to the Newtonian fluid when the time constant is zero ($\lambda$ = 0). Table 2 lists the parameters used for the CarreauYasuda model for blood in this study.
@@
    The recently developed viscosity model by Apostolides and Beris [28] is based on the Casson viscosity model. The parameters are functions of the constituents of blood like hematocrit and fibrinogen, whose values are adjusted using various flow-related data available in the literature. The apparent fluid viscosity is given by the following equation:
@@
    A threshold value of hematocrit is required to calculate the yield stress; this threshold is a quadratic function of the fibrinogen concentration $c_{f}$.
@@
    Table 3: Parameters considered for the Apostolidis and Beris blood viscosity model
    Blood Parameters for Apostolidis and Beris Viscosity Model Value
@@ 3 Flow distribution in a descending aorta
3.1 Grid Independence Test (GIT) and Validation
Simulations of the flow distribution in the descending aorta in patient-specific (Level 1) and its simplified versions (Levels 2 and 3) are performed using the finite volume method within Ansys-Fluent, a commercial flow solver. For the grid independence test for the three geometries, the steady flow problem is solved on four refined meshes (M1, M2, M3, and M4). The solver used in this study is also validated with the results of Nagargoje et al. [30] for a symmetric bifurcation with pulsatile inlet boundary conditions. The pulsatile waveform and the geometry of interest are shown in Figure 4. In Figure 5(a), the average velocity obtained is shown as a function of the inverse of the mesh element size. Following grid independence, the M3 mesh is selected for all further simulations. Validation against the work of [30]is shown in Figure 5(b). The $z$-component velocity is plotted on the centerline near the bifurcation for three-time instants of the oscillatory inflow condition. The validation results are seen to be satisfactory, which provides a check on the correctness of the solver.
@@ 3.2 Velocity
The systolic phase of the waveform (Figure 2) includes three key instants: $t_{1}$ = 0.2 s in the initial instant of increase in inflow velocity, $t_{2}$ = 0.3 s is the peak, and $t_{3}$= 0.45 s is the deceleration phase. The diastolic phase is analyzed at two time instants, $t_{4}$ = 0.6s and $t_{5}$ = 0.7s, representing flow fluctuation towards the end of the cycle. This section further describes the velocity characteristics over the cross-section in the form of centerline plots along with streamlines, contours, and vectors for the selected viscosity models and geometric simplification levels. In Figures 6 and 7, the distribution of the axial velocity component (denoted by $u$) is shown at the centerline of the cross-section located at a distance of $z=16D$ from the extended inlet as a function of the normalized $y$-coordinate. We show a comparison across two factors; in the first set of figures (Figure 6), we focus on comparing the viscosity model for a particular level of simplification, which shows subtleties in model predictions. In Figure7, we show a comparison across the level of simplification for any given viscosity model. This completes the picture and shows how modeling aspects are coupled with geometry details. In Figure 6, each level of geometry simplification is shown in a column, while rows represent distinct phases of the pulsatile flow cycle. The velocity profiles, normalized across the channel width, are depicted by red, blue, and black lines for the Newtonian, Carreau-Yasuda, and Apostolidis-Beris models, respectively. The axial velocity component distribution is shown at the selected time instants for the individual viscosity models for a particular geometry simplification level. An asymmetric velocity distribution is observed for Level 1 only. This is expected since the perturbation and undulations across the circumference at any axial position are retained in Level 1, which leads to an asymmetric flow in succeeding cross-sections. Interestingly, in Level 1, the flow profiles are nearly identical for all phases of the cycle for the Newtonian and Carreau-Yasuda viscosity models. This similarity is expected since the Newtonian and Carreau-Yasuda models predict similar viscosities at high shear rates (Figure 3). The velocity predictions for the Apostolidis and Beris viscosity model are slightly larger than the other two for all the phases. For Level 2 (Figure 6(b)), the geometry is simplified and made symmetric around the circumference, and symmetric flow profiles are observed for all the viscosity models at every phase of the pulsatile flow. Interestingly, at this level, the two non-Newtonian models - Carreau-Yasuda and Apostolidis and Beris- predict similar velocity profiles, whereas the Newtonian predictions show an overshoot near the centerline. Note that such differences at the centerline are known between Newtonian and non-Newtonian models for flows in a smooth duct where the latter needs a resolution of the shear rate singularity within the definition of apparent viscosity. For the simplest Level 3 in Figure 6(c) which is a tapered smooth tube, the predictions across viscosity models are largely symmetric and similar to each other, owing to the symmetry of the geometry.
@@
Next, in Figure 7, we show the velocity comparison across geometry simplifications for each viscosity model. The data shows an accuracy check for such simplifications across different viscosity models considered in the present study. Columns (a), (b), and (c) depict the velocity profiles for the Newtonian, Carreau-Yasuda, and Apostolidis and Beris viscosity models, respectively. As mentioned earlier, the rows represent distinct phases of the pulsatile flow cycle. Largely, the trends agree with each other, except for the asymmetry that is only visible for the patient-specific geometry (Level 1). The largest differences are observed at the start of the cycle (first row) and during the deceleration phase (third row). Interestingly, for the deceleration phase, only the predictions with the combination of Level 3 and Newtonian fluid are significantly more significant than the others. Thus, with respect to velocity profiles, the simplified geometry can capture the trends quite well for all viscosity models, except for the expected asymmetry in Level 1.
@@
These regions are particularly pronounced in areas with irregular geometry or abrupt changes in cross-section and surface perturbations. This backflow disrupts the unidirectional flow and interacts with the surrounding fluid, leading to flow separation and the formation of recirculation zones. The interplay between backflow, flow circulation, and pressure gradients indicates the sensitivity of the flow field to boundary conditions and geometric variations and reinforces the importance of retaining details in the geometry. Thus, a prediction accuracy loss is expected with geometry simplifications, some of which are examined next in Figure 9.

### v3

# Effects of Geometric Modelling and Blood Rheology in Patient-Specific Arterial Blood Flow Simulations with Speed-Accuracy Trade-Off Analysis

@@ Abstract
This study investigates the effects of geometry modeling and reduction on predictions of blood flow in patient-specific descending aorta obtained from Computed Tomography (CT) scan data, followed by speed-accuracy trade-off analysis using 3D simulations. Additionally, we show how wall shear stresses (WSS) can be estimated for such realistic arteries using cheaper simulations of highly idealized equivalent geometries. The study uses realistic inflow (with mean Re = 313.6 and peak Re = 1687 based on Newtonian viscosity) and pressure outlet boundary conditions with Newtonian and non-Newtonian blood viscosity models, including the one developed by Apostolidis and Beris [Alex J Apostolidis and Antony N Beris.”Modeling of the blood rheology in steady-state shear flows”, Journal of Rheology, 58(3):607–633, (2014)]. Similar studies exist to predict the progression of atherosclerosis, clots, aneurysms and thrombus formation in arteries. However, these assume idealized flow geometries. In contrast, real arteries possess significant amounts of roughness, perturbations, and undulations. To understand their impact along with blood rheology, studies are performed with two levels of geometry reduction. Our results show the effects of such simplifications on flow field predictions in conjunction with blood rheology models. The first level of reduction loses the local asymmetry but is able to reasonably approximate various parameters, capturing patterns with the correct magnitude while showing significant computational speed-up. However, further simplification to an idealized smooth geometry loses all information about the vortex structures in the flow field. The flow recirculation zones near wall perturbations progressively weaken as the geometry is smoothed. Broadly, for blood rheology models, no significant differences are observed between the various types used in this study. Interestingly, for all models, the idealized smooth geometry, combined with area correction, yields WSS estimates that closely approximate those of the actual artery. In this study, we also present a detailed speed-accuracy trade-off analysis useful for future applications. The number of iterations to convergence becomes progressively lower with geometry reduction, with nearly four times speed-up obtained with the first level of simplification. Further, the computational time for the Newtonian model remains approximately an order of magnitude lower than the non-Newtonian counterparts.
@@ 1 Introduction
Cardiovascular diseases (CVD) remain one primary reason for human mortality [1] and have grown in importance after the COVID-19 pandemic. The progression of CVD is directly related to the geometric details of the arteries [2]. CVD is known to aggravate other conditions as well [3]. Over the last few decades, greater computing power has enabled detailed computational studies on blood flow in the human arteries [4, 5]. Researchers have been interested in the link between atherosclerosis and fluid flow and related parameters, including the wall shear stress (WSS) [6, 7, 8, 9]. Increasingly, studies have used patient-specific geometries with the aim of contributing to medical applications [10]. Similar studies have been reported to predict the progression of atherosclerosis [11, 12], clots [13], aortic aneurysm[14] and formation of thrombus [15]. Significant studies have investigated the effect of stenosis on blood flow [16, 17, 18, 19]. However, in all these studies, the flow geometries have been assumed to be quite regular. In contrast, human arteries are far from regular and possess a significant amount of undulations in diameter as well as in the flow direction. The geometrical variations of arteries have been linked to the onset of atherosclerosis [20, 21, 22], thereby justifying its importance in the broader aspect of human health.
@@
Despite the importance of irregularities in geometry, researchers have rarely focused on their impact on blood flow. Yi et al. [23] performed simulations to study the impact of surface roughness on arterial stenosis. Shukla et al.[24] studied the effect of surface roughness on pulsatile flow in a tube with a periodic pattern of roughness in the form of sinusoids while Owen et al. [25] studied the effect of surface roughness on blood flow in a tube. To the best of our knowledge, these three studies come closest to the studying the impact of roughness from a biological perspective for blood flow analysis. Among them, the studies by Yi et al. [23] and Shukla et al. [24] use Newtonian fluid models. Owen et al. [25] used a variety of models, including Newtonian and non-Newtonian ones, to study the effect of an uneven geometry. However, the geometry and setup they created in the study can be argued to be far removed from reality. In a tube, they made a ”rough” patch with surface irregularities that extended along the length and compared fluid flow results between the rough patch and an equivalent smooth patch on the other side. This geometrical setup has two significant problems. Firstly, the roughness is distributed all over the surface, and a tube with a rough and smooth patch is not close to the surface features of arteries in our body. Secondly, the surface features comprise not just roughness but also changes on the scale of the artery diameter. Hence, irregularities also exist in the shape and size fluctuations along the axial length of the arterial tube. The flowing fluid experiences the combined effect of these irregularities. As discussed later, the real artery may deviate significantly from the circular shape over a given plane. Additionally, the equivalent diameter can also show variations along the axis. To the best of our knowledge, no previous study has investigated the effect of such shape and size perturbations, in addition to surface roughness, which is aimed for in the present study. Also, we use a variety of viscosity models to examine the sensitivity of the non-Newtonian nature of blood with geometric perturbations.
@@
One additional aspect studied here is the change obtained by the choice of the viscosity model. Calculations in these three equivalent geometric representations provide intricate details of the effects of geometric perturbations on the flow of blood and their connection with the chosen rheological model, apart from the effect of these details on the computational speed. Additionally, we also ask an inverse question. Since highly idealized geometries are computationally cheaper, is it possible to use such simulations to obtain an estimate of relevant haemodynamic parameters in the patient-specific, real geometry with all perturbations and undulations? In this study, we show the feasibility of this aspect for estimating WSS, which is extremely relevant in medical fields. The close agreement of such estimates and simulation results on the actual geometry, particularly for non-Newtonian models, opens up further possibilities for future applications.
The outcomes of the simulations are shown in terms of vortical patterns and hemodynamic parameters such as the time average wall shear stress (TAWSS) and oscillatory shear index (OSI). The accuracy losses are shown in terms of correlations defined in the article, which provides a numerical estimate of the loss of prediction accuracy across various simplification levels. The differences in flow patterns among the three levels are discussed in detail, with an understanding of the substantial reduction in CPU time. The results presented in this study may enable the simplification of complex arterial networks to increase the feasibility of meaningful full-body simulations in real-time.
@@ 2.1 Geometry
Figure 1: Level of geometrical aspects is shown, in (i) Level 1 of descending aorta extracted from CT-scan data, (i) Slice-based geometry of descending aorta (Level 1 is divided into 25 slices for the construction of level 2). (iii) Level 3 geometry based on the inlet and outlet diameter of Level 1 geometry. (iv) Cross-section area variation over the normalized z-axis for each level of simplification of the descending aorta.
@@ 2.2 Governing equations
Patient-specific and simplified models of the arteries are used to study the hemodynamics of blood flow under pulsatile flow conditions. In this respect, the incompressible and unsteady form of the continuity and momentum equations are solved in three dimensions with realistic boundary conditions with appropriate viscosity models. The incompressibility constraint in the form of the continuity equation and the applicable momentum equations for a variable viscosity fluid are given below.
@@ 2.2.1 Boundary Conditions
To solve the governing equations in the descending aorta and its simplified geometric versions, the following boundary conditions are applied:
@@
    Inlet conditions: A uniform velocity profile is imposed at the inlets of all geometries, combined with a pulsatile waveform to replicate the periodic nature of blood flow under physiological conditions. This pulsatile waveform captures the unsteady characteristics of flow and ensures a realistic simulation of the hemodynamic environment. The pulsatile velocity profile is derived from physiological measurements reported by Mills et al. [27]. The derived profile is fitted by a Fourier series of up to 18 terms to apply the inlet boundary condition with a mean Re = 313 and peak Re = 1687. To fully allow the inflow to develop into a Womersley profile, the inlet region of the artery is extended by a length of $10D$, where $D$ represents the characteristic diameter of the artery. This extension minimizes upstream flow irregularities owing to flow development and ensures a realistic inflow condition in the computational domain of interest. Similarly, an extension of $5D$ is added to the outlet to mitigate numerical reflections of an outflow boundary condition and maintain a smooth flow transition at the exit plane.
@@
    Outlet conditions: Realistic pulsatile gauge pressure is applied at the outlets of the geometries, as depicted in Figure 2. This boundary condition mirrors the dynamic pressure variation observed in the cardiovascular system, which are critical for accurately capturing the pressure-driven nature of blood flow. The pulsatile pressure profile is derived from physiological measurements carried out by Mills et al. [27], ensuring that they impose real-world conditions. These outlet conditions play a key role in maintaining the accuracy of flow characteristics, especially in scenarios where downstream vascular resistance and compliance significantly affect the upstream hemodynamics.
@@
    Wall conditions: The arterial walls in all geometries are assumed to be rigid, i.e. they do not deform under the influence of pressure and shear loading. A no-slip boundary condition is enforced at the walls, requiring the velocity at the wall to be zero. This assumption is widely used in computational blood flow simulations. Since a rigid wall is expected to display the strongest vortical structures.
@@ 2.2.2 Blood viscosity models
    In this context, $\mu_{0}$ and $\mu_{\infty}$ are the saturation values of viscosity at low and high shear rates, respectively. The parameter $P$ is known as the Yasuda exponent, with both $P$ and $n$ determining the degree of nonlinearity. The time constant $\lambda$ regulates the range of strain rates at which the transition to a power-law type regime is obtained. In this way, $\lambda$ is related to the relaxation time of the fluid. The model simplifies to the Newtonian fluid when the time constant is zero ($\lambda$ = 0). Table 2 lists the parameters used for the Carreau-Yasuda model for blood in this study.
@@
    The recently developed viscosity model by Apostolides and Beris [30] is based on the Casson viscosity model. The parameters are functions of the constituents of blood such as hematocrit concentration $\phi$ and fibrinogen cf, whose values are adjusted using various flow-related data available in the literature. The apparent fluid viscosity is given by the following equation:
@@
    A threshold value of hematocrit is required to calculate the yield stress; this threshold is a quadratic function of the fibrinogen concentration.
@@
    Table 3: Parameters considered for the Apostolidis and Beris blood viscosity model[30]
    Blood Parameters for the Apostolidis and Beris Viscosity Model Value
@@ 3 Flow distribution in a descending aorta
3.1 Grid Independence Test and Validation
3D simulations of the flow distribution in the descending aorta in patient-specific (Level 1) and its simplified versions (Levels 2 and 3) are performed using the finite volume method within ANSYS-Fluent®, a commercial flow solver. For the grid independence test for the three geometries, the steady flow problem is solved on four refined meshes (M1, M2, M3, and M4). The solver used in this study is also validated with the results of Nagargoje et al. [32] for a symmetric bifurcation with pulsatile inlet boundary conditions. The pulsatile waveform and the geometry of interest are shown in Figure 4. In Figure 5(a), the average velocity obtained is shown as a function of the inverse of the mesh element size. Following grid independence, the M3 mesh is selected for further simulations. Validation against the work of [32] is shown in Figure 5(b). The $z$-component velocity is plotted on the centerline near the bifurcation for three-time instants of the oscillatory inflow condition. The validation results are seen to be satisfactory, which provides a check on the correctness of the solver.
@@ 3.2 Velocity Profiles
The systolic phase of the waveform (Figure 2) includes three key instants: $t_{1}$ = 0.2 s in the initial instant of increase in inflow velocity, $t_{2}$ = 0.3 s is the peak, and $t_{3}$= 0.45 s is the deceleration phase. The diastolic phase is analyzed at two time instants, $t_{4}$ = 0.6s and $t_{5}$ = 0.7s, representing flow fluctuation towards the end of the cycle. This section further describes the velocity characteristics over the cross-section in the form of centerline plots along with streamlines, contours, and vectors for the selected viscosity models and geometric simplification levels. In Figures 6 and 7, the distribution of the axial velocity component (denoted by $u$) is shown at the centerline of the cross-section located at a distance of $z=16D$ from the extended inlet as a function of the normalized $y$-coordinate. We show a comparison across two factors; in the first set of figures (Figure 6), we focus on comparing the viscosity models for a particular level of simplification, which shows subtleties in model predictions. In Figure 7, we show a comparison across the level of simplification for any given viscosity model. This completes the picture and shows how modeling aspects are coupled with geometry details. In Figure 6, each level of geometry simplification is shown in a column, while rows represent distinct phases of the pulsatile flow cycle. The velocity profiles, normalized across the channel width, are depicted by red, blue, and black lines for the Newtonian, Carreau-Yasuda, and Apostolidis-Beris models, respectively. The axial velocity component distribution is shown at the selected time instants for the individual viscosity models for a particular geometry simplification level. An asymmetric velocity distribution is observed for Level 1 only. This is expected since the perturbation and undulations across the circumference at any axial position are retained in Level 1, which leads to an asymmetric flow in succeeding cross-sections. Interestingly, in Level 1, the flow profiles are nearly identical for all phases of the cycle for the Newtonian and Carreau-Yasuda viscosity models. This similarity is expected since the Newtonian and Carreau-Yasuda models predict similar viscosities at high shear rates (Figure 3). The velocity predictions for the Apostolidis and Beris viscosity model are slightly larger than the other two for all the phases. For Level 2 (Figure 6(b)), the geometry is simplified and made symmetric around the circumference, and symmetric flow profiles are observed for all the viscosity models at every phase of the pulsatile flow. Interestingly, at this level, the two non-Newtonian models - Carreau-Yasuda and Apostolidis and Beris- predict similar velocity profiles, whereas the Newtonian predictions show an overshoot near the centerline. Note that such differences at the centerline are known between Newtonian and non-Newtonian models for flows in a smooth duct where the latter needs a resolution of the shear rate singularity within the definition of apparent viscosity. For the simplest Level 3 in Figure 6(c) which is a tapered smooth tube, the predictions across viscosity models are largely symmetric and similar to each other, owing to the symmetry of the geometry.
@@
Next, in Figure 7, we show the velocity comparison across geometry simplifications for each viscosity model. The data shows an accuracy check for such simplifications across different viscosity models considered in the present study. Columns (a), (b), and (c) depict the velocity profiles for the Newtonian, Carreau-Yasuda, and Apostolidis and Beris viscosity models, respectively. As mentioned earlier, the rows represent distinct phases of the pulsatile flow cycle. Largely, the trends agree with each other, except for the asymmetry that is only visible for the patient-specific geometry (Level 1). The largest differences are observed at the start of the cycle (first row) and during the deceleration phase (third row). Interestingly, for the deceleration phase, only the predictions with the combination of Level 3 and Newtonian fluid are significantly more than the others. Thus, with respect to velocity profiles, the simplified geometry can capture the trends quite well for all viscosity models, except for the expected asymmetry in Level 1.
@@
These regions are particularly pronounced in areas with irregular geometry or abrupt changes in cross-section and surface perturbations. The backflow generated disrupts the unidirectional flow and interacts with the surrounding fluid, leading to flow separation and the formation of recirculation zones. The interplay between backflow, flow circulation, and pressure gradients indicates the sensitivity of the flow field to boundary conditions and geometric variations and reinforces the importance of retaining details in the geometry. Thus, a prediction accuracy loss is expected with geometry simplifications, some of which are examined next in Figure 9.